OpenStudy (anonymous):

A college graduate enlisted in the army for 4 years. So far, he has served 9 months. What percent of his enlistment has he served to the nearest whole number ?

OpenStudy (anonymous):

\[\frac{9}{4*12}*100=\frac{75}{4}=18.75\% \]

OpenStudy (anonymous):

THX!!!

OpenStudy (anonymous):

You're welcome.